# Robots that Suggest Safe Alternatives

### v1

@@ V-B What is the Benefit of SALT as a Runtime Monitor?
Figure 5: Figure 5: User Alignment Success Across 1000 Initial Conditions. (Left) Alternative alignment in the navigation setting, where $g_{\mathcal{H}}^{1}=(3,-3),g_{\mathcal{H}}^{2}=(3,0),g_{\mathcal{H}}^{3}=(3,3)$. (Right) Alternative alignment in manipulation. Although the navigation has high alignment, people did not deem any other objects acceptable when $g_{\mathcal{H}}$ = RedMug.
V-C How Acceptable are SALT’s Proposed Alternatives?
@@
Alternatives Dataset. To measure how acceptable are SALT’s alternatives, we obtained a validation dataset of initial goal ($g_{\mathcal{H}}$) and acceptable alternative pairs annotated by 10 expert users from labs at Carnegie Mellon and UC San Diego. In the navigation scenario, people were shown three initial goals at the top, center, and bottom of the corridor. They were asked the maximum acceptable distance an alternative goal could be from the original goal. For manipulation, people were shown a tabletop with three objects (as in Fig. 1), and asked which alternative objects would be acceptable given an initial object.
@@
Approach. For each initial goal $g_{\mathcal{H}}$ in the validation dataset, we queried SALT to obtain our algorithm’s suggested alternatives. We perform this across a suite of initial robot states $s$. In both environments, we queried 1,000 random initial conditions for each goal and saved our algorithm’s suggestion, $g_{\mathcal{R}}$. Note that SALT will return the initial proposed goal (i.e., $g_{\mathcal{R}}=g_{\mathcal{H}}$) if it is safe and live.
@@
Metrics. We measure alternative alignment (%): given an initial goal $g_{\mathcal{H}}$, if the alternative goal proposed by SALT matches or if the input goal is safe and live, then alignment is a success. We compute mean and standard deviation across all users and initial $s$, and report results per initial goal.
Results: Quantitative. Results are shown in Figure 5. In navigation, five out of ten users wanted a bound less than $1m$ for $g_{\mathcal{H}}^{3}=(3,3)$ and $g_{\mathcal{H}}^{1}=(3,-3)$ (top and bottom goals), while four users wanted a bound less than $1m$ for $g_{\mathcal{H}}^{2}=(3,0)$ (middle goal in Figure 3, left). Despite this strict bounding, SALT was able to successfully propose an average aligned alternative of above 90% for all three goals across all users for 1,000 random initial conditions. In manipulation, when $g_{\mathcal{H}}=\text{RedBowl}$ or $g_{\mathcal{H}}=\text{BrownBowl}$, 80% of users preferred the other bowl as an alternative object. Figure 5 shows that SALT is able to reliably choose alternatives that align for bowl goals (93.84% and 100% respectively average across users and initial conditions). When $g_{\mathcal{H}}=\text{RedMug}$, 80% of users wanted neither of the alternative objects; thus, successful alignment between SALT and a user happened 55.85% on average across users and initial conditions (right, Figure 5). We hypothesize this is because of the semantic difference between the original goal and the suggestion: for many purposes, a bowl won’t do as a replacement for a mug.
@@
Results: Qualitative. Figure 3 center shows an out-of-distribution human-requested goal. SALT realizes that rolling out the policy on that goal would lead to collision, and proposes a nearby alternative in the center of the room that it can successfully execut (right, Figure 3). Figure 1 shows our algorithm in manipulation where a user first asks for the red bowl to be picked up. Realizing the it would likely mis-grasp the red bowl and fail, SALT proposes to pick up the brown bowl (right, Figure 1), and safely completes the task.
@@ VI Conclusion
In this work, we propose SALT: a framework for robots to monitor if their goal-conditioned policies can safely accomplish a given goal, and automatically suggest safe alternatives when they cannot. By actively proposing a safe alternative pre-execution, we can not only minimize possible safety violations, but also human intervention efforts. For safety, we find that open loop monitors that interpret safety as uncertainty cannot differentiate clearly between a safe and unsafe state, while our closed-loop monitor reliably detects predicts success and failure. In future work, we seek to investigate goal representations that are obtained via language descriptions or images of goal states, and robustify the reach-avoid value network by injecting a disturbance.

### v2

@@ V-B What is the Benefit of SALT as a Runtime Monitor?
V-C How Should SALT Reason About “Similar” Alternatives?
We hypothesize that suggesting similar alternatives may require a semantic representation of the goals—capturing visual or functional properties of the object—and a corresponding similarity measure. In this section, we ablate the encoder models $\mathcal{E}(\cdot)$ and similarity measure $d(\cdot)$ used within our SALT framework. We also study how similar alternatives change as a function of human intent $\theta$ in Equation III.
Methods: Encoders, Intents, and Similarity Measures. We investigate three methods that leverage different goal representations and similarity measures: CosineSim, LLM, and SIRL. In our experiments, CosineSim uses a textual description of the goal (e.g., as returned by a semantic object detector) and measures similarity via the cosine similarity between the textual embedding of the human’s original goal ($g_{\mathcal{H}})$ and any alternative goal ($g$). We use the pre-trained BERT [42] sentence-transformer model to obtain the textual embedding. Mathematically, given a language description $\mathcal{L}_{g}$ of a goal $g$ and a textual description of the intent $\mathcal{L}_{\theta}$, our encoder produces an embedding vector $\mathcal{E}_{\text{BERT}}(\mathcal{L}_{g};\mathcal{L}_{\theta})=\vec{w}_{g}$ and our similarity measure is $d:=\vec{w}_{g}\cdot\vec{w}_{g_{\mathcal{H}}}/||\vec{w}_{g}||~{}||\vec{w}_{g_{\mathcal{H}}}||$. LLM Fuzzy Matching [43] also uses a textual representation of the goal and intent, but uses a pre-trained large language model (LLM) to directly reason about the semantic similarity (without looking at the embedding similarity). In our experiments, we use GPT-4o [44] as our language model. Mathematically, $d:={\color[rgb]{0.0703125,0.69921875,0.921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.69921875,0.921875}\textbf{LLM}}(\mathcal{L}_{g},\mathcal{L}_{g_{\mathcal{H}}};\mathcal{L}_{\theta},\mathcal{P})$, where the prompt to the LLM is $\mathcal{P}=$ “The user intends to $\mathcal{L}_{\theta}$. Given $\mathcal{L}_{g_{\mathcal{H}}}$, which item is the closest related to it?”.
Unlike LLM and CosineSim (which use pre-trained language models), SIRL requires training a personalized representation which explicitly learns an end-user’s notion of similarity from their preference data, enabling us to study how a personalized model of similarity influences our SALT framework compared to pre-trained models that are not fine-tuned on individual data. This model uses a privileged, hand-engineered feature space $\Phi_{g}\in\mathbb{R}^{m}$ as input based on any given goal object $g$; for example, in our experiments, given a $g$ is a red cup, $\Phi_{g}$ would be a 20-dimensional vector consisting of the object’s RGB color values and functional and material properties. We train an intent-parameterized encoder $\mathcal{E}_{{\color[rgb]{0.80078125,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.80078125,0,0.3984375}\textbf{SIRL}}}(\Phi(g);\theta)=\phi^{\theta}_{g}$ via contrastive learning, which returns an embedding vector $\phi^{\theta}_{g}\in\mathbb{R}^{n},~{}n<m$ that represents the most relevant features of a goal given the user’s intent. Finally, SIRL measures similarity via L2 distance in embedding space: $d:=||\phi^{\theta}_{g}-\phi^{\theta}_{g_{\mathcal{H}}}||_{2}$.
Figure 5: Figure 5: Accuracy of Different Similarity Measures. (Left Three Plots) RelRank scores for 5 aligned $g_{\mathcal{H}}$ are scored across baselines and intents. (Right Plot) TopRank scores for 5 aligned $g_{\mathcal{H}}$ are scored across baselines and intents.
Implementation Details. SIRL learns relevant similarity features by asking the end user to select the two most similar goals given a triplet of goals. Throughout this section, we use simulated human data for training SIRL and evaluation, enabling us to have access to a ground-truth representation of the human’s notion of similarity given the intent. SIRL is trained on triplets $\mathcal{D}_{\theta}=\{(\Phi^{i}_{g_{1}},\Phi^{i}_{g_{2}},\Phi^{i}_{g_{3}})\}^{K}_{i=1}$, which are feature spaces corresponding to three distinct goals in the environment given a human intent $\theta$. The simulated human ranks the two most similar ones using their ground-truth intent-relevant features. We train the encoder $\mathcal{E}_{{\color[rgb]{0.80078125,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.80078125,0,0.3984375}\textbf{SIRL}}}(\Phi(g);\theta)=\phi^{\theta}_{g}$ to minimize the loss from [45]:
$\mathcal{L}(\phi^{\theta}_{g})=\sum_{i=1}^{|\mathcal{D}_{\theta}|}\mathcal{L}_{trip}(\Phi_{g}^{i},\Phi_{g_{+}}^{i},\Phi^{i}_{g_{-}})+\mathcal{L}_{trip}(\Phi_{g_{+}}^{i},\Phi_{g}^{i},\Phi^{i}_{g_{-}})$
where $\Phi_{g}^{i}$ and $\Phi_{g_{+}}^{i}$ are ranked as most similar and $\Phi_{g_{-}}^{i}$ is most dissimilar. $\mathcal{L}_{trip}(\Phi_{g}^{i},\Phi_{g_{+}}^{i},\Phi^{i}_{g_{-}})$ is the triplet loss [46] that uses $\Phi_{g}^{i}$ as the anchor, $\Phi_{g_{+}}^{i}$ as the similar example and $\Phi^{i}_{g_{-}}$ as the dissimilar example. This loss function pushes together embeddings for similar objects with respect to the intent-relevant features while pushing apart embeddings for dissimilar objects.
Evaluation Setup. We use 10 kitchen objects (e.g., cups, bowls, mugs, pitchers, teapots, ramekins) from the Google Research dataset [47] and study three increasingly complicated user intents: $\theta_{1}=$ Sorting Kitchen by the Same Color, $\theta_{2}=$ Microwaving Soup, and $\theta_{3}=$ Serving Wine at Formal Dinner. For each intent, we select 5 initial goal objects from the dataset that make sense for the intent. Given one of these five objects, we query each method for its similarity score compared to all other 9 objects. We determine the ground-truth similarity score using a simulated human with privileged knowledge about their intent-relevant features.
Metrics. We want to measure how well the distance functions in each method captures the relative distance between any pairs of objects. We use two metrics: TopRank returns one if the most similar item outputted by a similarity measure is the most similar item for the simulated human (and zero otherwise); RelRank returns a real-valued score (between 0 to 1) which measures how correctly a method ranks all other goals relative to a given goal [^2] Our metric is inspired by from Kendall’s rank correlation coefficient [48]. However, while Kendall’s penalizes (with -1) for every incorrectly ordered pair, our metric does not (assigns 0). This modification is appropriate for our context, where the top rankings are of more practical importance (since we want to return maximally similar alternatives). Thus, we should not punish the method for having noisy bottom rankings since the differences may not represent meaningful distinctions. . Mathematically, let $g_{\mathcal{H}}$ be a given goal and let $\mathcal{G}^{*}_{g_{\mathcal{H}}}=\{g_{1},g_{2},...g_{n}\}$ be a list of all other $n$ goals ranked by how similar they are to $g_{\mathcal{H}}$ by the simulated human. We define $\mathcal{G}^{\omega}_{g_{\mathcal{H}}}$ to be a list of all the $n$ goals ranked by their similarity to $g_{\mathcal{H}}$ by any method, $\omega\in\{{\color[rgb]{0.80078125,0,0.3984375}\definecolor[named]{pgfstrokecolor}{rgb}{0.80078125,0,0.3984375}\textbf{SIRL}},{\color[rgb]{0.13671875,0.44921875,0.36328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.13671875,0.44921875,0.36328125}\textbf{CosineSim}},{\color[rgb]{0.0703125,0.69921875,0.921875}\definecolor[named]{pgfstrokecolor}{rgb}{0.0703125,0.69921875,0.921875}\textbf{LLM}}\}$. Let $r_{\mathcal{G}^{*}_{g_{\mathcal{H}}}}(g_{i})$ be the rank position of goal $g_{i}$ in the ordered list $\mathcal{G}^{*}_{g_{\mathcal{H}}}$ and let $R_{\mathcal{G}^{*}_{g_{\mathcal{H}}}}(g_{i},g_{j})=\text{sign}[r_{\mathcal{G}^{*}_{g_{\mathcal{H}}}}(g_{j})-r_{\mathcal{G}^{*}_{g_{\mathcal{H}}}}(g_{i})]$. Here, $R_{\mathcal{G}^{*}_{g_{\mathcal{H}}}}(g_{i},g_{j})$ returns 1 if $g_{i}$ is more similar to $g_{\mathcal{H}}$ than $g_{j}$ and -1 for the converse (our lists assume no ties). Finally, we define an indicator function $\mathsf{1}[R_{\mathcal{G}^{*}_{g_{\mathcal{H}}}}(g_{i},g_{j})=R_{\mathcal{G}^{\omega}_{g_{\mathcal{H}}}}(g_{i},g_{j})]$ that returns 1 if the two lists have the same relative rankings of the goals and 0 otherwise. The RelRank metric is then:
RelRank $(\mathcal{G}^{*}_{g_{\mathcal{H}}},\mathcal{G}^{\omega}_{g_{\mathcal{H}}})=$ (6)
$\frac{1}{\binom{n}{2}}$ $\sum_{1\leq i<j\leq n}\mathsf{1}[R_{\mathcal{G}^{*}_{g_{\mathcal{H}}}}(g_{i},g_{j})=R_{\mathcal{G}^{\omega}_{g_{\mathcal{H}}}}(g_{i},g_{j})].$
Results: Most Similar Goal Accuracy. In the right of Figure 5 we show the TopRank results averaged across all five goals per each intent. SIRL most consistently ranks the human’s preferred goal as most similar (in total, 12) compared to CosineSim and LLM (9 for both). Furthermore, the explicit training of SIRL makes it more robust to intents while LLM and CosineSim (parameterized by the intent) sometimes ignore critical intent features. (e.g., for $\theta_{1}=$Sorting Kitchen by Color and $g_{\mathcal{H}}=\text{RedMug}$, both methods return White Mug, neglecting color). We hypothesize that this occurs because the models are only approximately optimal in their rankings and selecting the top-1 choice requires precision; for example, we observed that the ground-truth top choice typically appeared within the top-3 results of each method.
Results: Overall Similarity Measure Accuracy. We report the RelRank metric (which quantifies the overall performance of a similarity measure) in the three plots left of Figure 5 for each intent $\theta_{1},\theta_{2},\theta_{3}$. Across all intents and goals, SIRL’s similarity measure is more consistently accurate compared to CosineSim and LLM, and is the best performing measure for intent $\theta_{3}$. This is because SIRL is optimized to solve a personalized metric learning problem—identifying an embedding space that understands similarity according to the user’s internal state—while the other two approaches that use pre-trained models have an implicit semantic understanding of similarity. As the intents become increasingly complicated (e.g., $\theta_{3}=$Serving Wine, CosineSim and LLM’s pre-trained similarity struggle to capture how the human evaluates these alternatives, while SIRL maintains performance due to its privileged access to human’s internal states. Between the two pre-trained approaches, LLM typically out-performs CosineSim in terms of accuracy.
Our main takeaway is that personalized preference data enables more accurate goal representations and better similarity measures, particularly for complex intents. However, off-the-shelf language models and LLM Fuzzy Matching can still be valuable semantic similarity measures that require no additional training data and provide a more naturalistic and intuitive interface for people to specify their goals (e.g., via language rather than explicit featurizations of the world).
V-D How Acceptable are SALT’s Proposed Alternatives?
Finally, we study SALT’s overall performance at detecting when the robot can safely accomplish a desired goal and suggesting a similar safe alternatives when it cannot.
@@ V-D How Acceptable are SALT’s Proposed Alternatives?
Alternatives Dataset. To measure how acceptable are SALT’s alternatives, we obtained a validation dataset of initial goal ($g_{\mathcal{H}}$) and acceptable alternative pairs annotated by 20 expert users from labs at Carnegie Mellon and UC San Diego. In the manipulation setting, people were shown a tabletop with three objects (as in Fig. 1) with two distinctive intents: $\theta_{1}=$ Drinking Soup and $\theta_{2}=$ Sorting Kitchen by the Same Color. These users were asked which alternative objects would be acceptable given an initial object.
@@
Approach. For each initial goal $g_{\mathcal{H}}$ in the validation dataset, we queried SALT to obtain our algorithm’s suggested alternatives. We perform this across a suite of initial robot states $s$. We queried 1,000 random initial conditions for each goal and saved our algorithm’s suggestion, $g_{\mathcal{R}}$. Note that SALT returns the initial goal (i.e., $g_{\mathcal{R}}=g_{\mathcal{H}}$) if it is safe and live.
@@
Metrics. We measure alternative alignment (%): given an initial goal $g_{\mathcal{H}}$, if the alternative goal proposed by SALT matches or if the input goal is safe and live, then alignment is a success. We compute mean and standard deviation across all users and initial $s$, and report results per initial goal. We compare alignment across several distance measures: Euclidean, CosineSim, SIRL, and LLM.
Results: Quantitative. Alternative alignment results are shown in Figure 6. On aggregate, we find that SALT can detect if the original goal was safe and suggest safe alternatives that strongly align with human preferences (alignment scores between 70%-100% when using the SIRL or LLM objective functions. We note that when $\theta_{1}=$Drinking Soupand $g_{\mathcal{H}}=\text{RedMug}$, 35% of the users wanted neither alternative, dropping the maximum possible alignment rate.
Furthermore, we break down our results into only those scenarios where the human’s initial goal ($g_{\mathcal{H}}$) was not safe, and thus the robot had to suggest an alternative. The alignment scores for these scenarios are shown in Figure 6 in a cross-hatched pattern (called AltSuggest). Note that for $g_{\mathcal{H}}=\text{BrownBowl}$, the robot is always capable of grabbing this initial object safely, and thus it is not part of the AltSuggest breakdown. First, we see that Euclidean has low alignment for most of the AltSuggest scenarios, highlighting the need for semantic similarity measures. For intent $\theta_{1}=$Drinking Soup and $\theta_{2}=$Sorting by Color, CosineSim’s alignment rate drops significantly (to 2.13% and 10% respectively) when $g_{\mathcal{H}}=\text{RedBowl}$, indicating that this similarity measure is not capable of making the closed-loop system suggest safe and aligned alternatives. Consistent with Section V-C, we see that when SALT uses SIRL and LLM, the overall system is able to consistently return safe and similar alternatives that align with an end-user’s notion of similarity.
@@
Results: Qualitative. Figure 1 shows our algorithm in manipulation where a user first asks for the red bowl to be picked up with the intent $\theta_{1}=$ Drinking Soup. Realizing the it would likely mis-grasp the red bowl and fail, SALT proposes to pick up the brown bowl with SIRL (right, Figure 1), and safely completes the task.
Figure 6: Figure 6: User Alignment Success Across 1,000 Initial Conditions. Euclidean is in grey, CosineSim in green, SIRL in pink, and LLM in blue. Scenarios where the initial goal $g_{\mathcal{H}}$ was not safe and thus the robot had to suggest an alternative are denoted as AltSuggest. SIRL and LLM consistently out-perform Euclidean and CosineSim when it comes to guiding SALT to generate safe and aligned alternatives.
@@ VI Conclusion
In this work, we propose SALT: a framework for robots to monitor if their goal-conditioned policies can safely accomplish a given goal, and automatically suggest safe alternatives when they cannot. By actively proposing a safe alternative pre-execution, we can not only minimize possible safety violations, but also human intervention efforts. For safety, we find that open loop monitors that interpret safety as uncertainty cannot differentiate clearly between a safe and unsafe state, while our closed-loop monitor reliably predicts success and failure. One limitation of our work is that we rely on privileged state information about the environment as well as goal representations to quantify safety and suggest alternatives. Another limitation is the use of a simulated human during our similarity measure evaluation, since it may not accurately model real user preferences. We are excited about further verifying this with a rigorous user study. In future work, we seek to investigate image-based goal-conditioned policies, as well as image representations of goals and similarity queries. One exciting future work is reasoning about task-level alternatives, rather than high-level alternatives like goals.